OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 2x + 4y = 12. slope = 2; y intercept at (0, 12) slope = one half; y intercept at (0, 3) slope = one half; y intercept at (0, 4) slope = -one half; y intercept at (0, 3)

OpenStudy (anonymous):

its the first one

OpenStudy (anonymous):

thanks Katie for not just rushing out and say answer an leave would you know how to explain an little

OpenStudy (anonymous):

move twelve over so that the equation equals 0 2x+4y-12=0 subtract 4y from both sides2x-12=4y divide all by zero so the first answer is correct.1/2x-3=y slope is 1/2 and y intercept is -3

OpenStudy (anonymous):

I just need help